acconting

Suppose the following information is available for Callaway Golf Company for the years 2022 and 2021. (Dollars are in thousands, except share information.)



2022
2021

Net sales


1,129,000
1,136,000

Net income (loss)


84,462
70,965

Total assets


855,338
838,078





Share information



Shares outstanding at year-end


69,500,000
71,270,000

Preferred dividends


0
0


There were 78,130,000 shares outstanding at the end of 2020.

(a)

What was the company’s earnings per share for each year?
Answer #1

Answer

· Requirement [a]

2022

$1.20

2021

$0.95

· Working

2021

2022

A

Net Income (loss)

$70,965,000

$84,462,000

B

Preferred dividends

$0

$0

C = A - B

Net Income for common stockholders

$70,965,000

$84,462,000

D

Beginning shares outstanding

                78,130,000

             71,270,000

E

Ending shares outstanding

                71,270,000

             69,500,000

F = (D+E) / 2

Weighted average no of common stock shares outstanding

                74,700,000

             70,385,000

G = C/F

Earnings per share

$0.95

$1.20